# Float Self-Tagging

### v1

@@ 5.1. Tag Testing
The cost of the check does not depend on the number of tags tested. This can be advantageous if a tagged pointer is used for boxed floats, say with tag b. In that case, a check for a float (either tagged or boxed) can be done by adding b to the tag set. Once it is known that an object is a float, it is easy to check for the single tag b to discriminate between tagged and boxed representations.
@@
To convert a f64 to an object, it must first be converted to an i64, and then the high bits of the exponent must be aligned with the 3 low bits using a rotate instruction. If those 3 low bits are one of 000, 011, or 100 then the conversion is done. Otherwise, an out of line routine can be called to allocate a boxed float and return an appropriately tagged pointer. If the f64 value is in the 64-bit float register xmm0, then the following x86 code will set register rax to the corresponding object:
@@
rol rax,#4 # rotate left 4 places
@@
jc done
call allocate_boxed_float
@@
Assuming the initialization of the tag mask register is amortized over multiple float-related operations, the cost for the hot path is low: 3 register-to-register operations and an easily predictable conditional jump.
@@
To convert an object (that is known to be a float) to a f64, the tag must be tested to see if it is a tagged float. If it is, rotating the bits 4 places to the right will restore the bits of the f64 result. Otherwise, a memory read gets the f64 result out of the boxed float. If the object value is in the 64-bit register rax, then the following x86 code will set register xmm0 to the corresponding f64:
bt x,ax # assumes bx initialized to 0x1919 elsewhere
jc rotate
mov xmm0,[rax+offset] # xmm0 $\leftarrow$ value field of boxed float
@@
rotate:
ror rax,#4 # rotate right 4 places
@@
Here too the cost is low for the hot path: 3 register-to-register operations and an easily predictable conditional jump.
@@ 5.4. C Implementation
This section explains how the above operations can be implemented in C, as this is a common implementation language, in particular it is used by the Bigloo Scheme to C compiler used in the experiments of Sections 3 and 4.
Checking for a set of tags can be implemented with the following pure C function, which is easily inlined by a C compiler:
#define TAG_SET ((1<<4)|(1<<3)|(1<<0)) /* 0x19 for tags 100, 011, and 000 */
inline bool has_tag_4_or_3_or_0(int64_t n) {
return (((uint32_t)1 << n) & (~(uint32_t)0/0xff * TAG_SET)) != 0;
The expression (uint32_t)1<<n implicitly computes a shift count mod 32 and the bitwise-and therefore tests the bit at index n mod 32 of its second operand. The code uses the compile-time constant ~(uint32_t)0/0xff * TAG_SET that repeats the 8 bit mask TAG_SET 4 times to fill a 32 bit word, giving 0x19191919. This achieves the equivalent of testing the bit at index n mod 8 of TAG_SET.
Using this pure C definition with GCC version 13.2.0 and clang version 18.1.3 on x86-64 results in both cases in the non-destructive approach based on the bit test instruction. On older versions of those compilers that would implement this with a destructive shift instruction, the bit test approach can be achieved using the following definition that uses an asm statement:
inline bool has_tag_4_or_3_or_0(int64_t n) {
@@
Boxing and unboxing operations need to convert between a float value and its bit representation. This can be achieved portably with a union type whose fields are of type f64 and i64. Moreover, boxing and unboxing need to rotate the bit representation of the float. Although C does not provide an operator for bit rotation, both GCC and clang recognize an equivalent pair of shifts and generate a single machine rotate instruction. The implementation in C is:
@@
#define ROTR(n,s) ((int64_t)(((uint64_t)n >> s) | ((uint64_t)n << (64 - s))))
@@
int64_t result = ROTR(((union di)f).i, 60);
if (has_tag_4_or_3_or_0(result)) return result;
return allocate_boxed_float(f);
@@
int64_t result = ROTR(o, 4);
if (has_tag_4_or_3_or_0(result)) return ((union di)result).d;
return value_of_boxed_float(o);
@@ 6. Self-tagging on 32-bit architectures
Although previous sections only considered double-precision floats on 64-bit architectures, it is straightforward to adapt float self-tagging to single-precision floats on a 32-bit architecture. The core idea, which is to superimpose a tag to a sequence of bits likely to appear in practice, can be adapted to any machine word size.
On 32-bit implementations, it is frequent to have four 2-bit tags and heap-allocated objects aligned to 32-bit words. One option is to select two tags for float self-tagging, one tag for small integers, and assign the remaining tag to generic pointers. A similar interval analysis as in Section 4 leads to the choice of 01 and 10 which capture the interval from $2^{-63}$ (included) to $2^{65}$ (excluded), roughly $1.1\times 10^{-19}$ to $3.7\times 10^{19}$, and the equivalent negative interval illustrated in Figure 14.
Figure 14. Logarithmic scale number line of single-precision float values unboxed by self-tagging.
This tag selection covers half of all floats, but does not capture the values $\pm$0.0. However, the experiment from Section 4.1 demonstrated that, by preallocating $\pm$0.0, it can improve performance.
As a final remark, self-tagging can be applied in an implementation using two 1-bit tags with a single tag that captures half of all floats (provided the only free tag is not assigned to small integers). In this context, one may consider superimposing the tag with the float sign bit to capture all positive floats.
@@ 7. Related Work
Unboxing floats is a longstanding problem in dynamic language implementation. In the last decades, little progress has been made improving the encoding of floats, with most implementations either using boxed floats or suffering from the overhead of NaN-tagging on pointers. Thus, more general strategies that use data flow analysis (Henglein, 1992; Serrano and Feeley, 1996; Petersen and Glew, 2012; Lindahl and Sagonas, 2003) and partial evaluation (Würthinger et al., 2017) have been developed to find locations where the type of a value is known across its lifecycle. This allows the generation of specialized code that safely handles fully unboxed, untagged values. Such strategies are not specific to unboxing floats. Rather, they tackle the more general problem of inferring types in dynamic languages, which allows dropping type information in either tagged pointers or NaN-boxed pointers. While more limited in scope, self-tagging solves the problem of boxed floats with a new encoding that involves no program analysis.
@@
Self-tagging is straightforward to add to compilers that already use object tagging. Therefore, implementations that represent floats as tagged pointers could benefit from it with minimal implementation effort. Such popular implementations include CPython (Gross, 2023), CRuby (Shaughnessy, 2013) and Google’s V8 (Sheludko and Solanes, 2020). Compilers that use NaN-tagging, such as Mozilla’s SpiderMonkey (Foundation, 2024) and Apples’s JSC (Inc., 2024) can also benefit from the lower impact that self-tagging has on the performance of non-float types compared to NaN-tagging.
@@
Due to the shortcomings of tagged pointers and NaN-tagging, additional strategies are used to avoid boxing. A straightforward approach is to provide homogeneous data structures such as TypedArray in JavaScript (International, 2015) and numpy’s float64 arrays in Python (Harris et al., 2020). Type homogeneity then allows unboxing values within the collection. More generally, compilers can detect data homogeneity at run time and use context-dependent storage strategies to store data in a way that prevents pointer chasing (Bolz et al., 2013). Since self-tagging unboxes all floats in practice, it effectively prevents such pointer chasing in the case of floats without static of run time analysis.

### v2

@@ 5.2. Multiple Tag Testing
The cost of the check does not depend on the number of tags tested. This can be advantageous if a tagged pointer is used for heap-allocated floats, say with tag h. In that case, a check for a float (either self-tagged or heap-allocated) can be done by adding h to the tag set. Once it is known that an object is a float, it is easy to check for the single tag h to discriminate between the self-tagged and heap-allocated representations.
@@
To convert a f64 to an object, it must first be converted to an i64, and then a constant $bias$ must be added modulo $2^{64}$ followed by a bit-rotation by $r$ places left. For the 1-tag variant, $bias=(1+2\times tag)\times 2^{58}$ and $r=5$. For the 2-tag, 3-tag, and 4-tag variants, $bias=0$ (no bias) and $r=4$. After this there is a check to verify if the resulting object’s tag is in the set of self-tagged tags. If this is the case the conversion is done, otherwise an out of line routine can be called to heap-allocate a float and return an appropriately tagged pointer. As an example, if the f64 value is in the 64-bit float register xmm0, then the following x86 code will set register rax to the corresponding object when the 3-tag self-tagging variant is used:
@@
rol rax, 4 # rotate left 4 places
@@
jnc heap_alloc_float # tag is not one of 000, 011, or 100?
@@
Assuming the initialization of the tag mask register (bx) is amortized over multiple float-related operations, the cost for the hot path is low: 3 register-to-register operations and an easily predictable conditional jump.
The following x86 code is appropriate for the 1-tag self-tagging variant using 000 as the tag:
mov rax, xmm0 # rax $\leftarrow$ xmm0
add rax, rbx # assumes rbx initialized to 1<<58 elsewhere
rol rax, 5 # rotate left 5 places
test al, 7
jnz heap_alloc_float # tag is not 000?
done:
Here also, the cost for the hot path is just a bit more due to the bias: 4 register-to-register operations and an easily predictable conditional jump.
For the variant explained in Section 4.4 which uses 2 tags with no special handling of zeroes, it is more efficient to do the rotation before adding the bias because then the bias is the small constant $tag_{1}$, and in the special case of $tag_{1}$=000 the bias vanishes. For example, here is the code when $tag_{1}$=011:
mov rax, xmm0 # rax $\leftarrow$ xmm0
rol rax, 5 # rotate left 5 places
add al, 3 # sufficient to modify the lower 8 bits of rax
bt bx, ax # assumes bx initialized to 0x0c0c elsewhere
jnc heap_alloc_float # tag is not one of 010 or 011?
done:
The add instruction uses a small constant as a bias, contrary to the code for the 1-tag version (large constants typically need to be setup in a register, which is additional work even if it is amortized). Moreover the 1-tag version cannot eliminate the bias, but here the add instruction can be removed when $tag_{1}$=000 which makes it slighly faster than the 1-tag variant and as efficient as the 3-tag variant: 3 register-to-register operations and an easily predictable conditional jump.
@@
To convert an object (that is known to be a float) to a f64, the tag must be tested to see if it is a self-tagged float. If it is, the boxing operations are done in reverse order and inverted, replacing the rol by a ror (rotate-right), and if there is a bias, replacing the add by a sub (subtract). If it is not a self-tagged float, a memory read gets the f64 result out of the heap allocated float. As an example using the 3-tag variant, if the object value is in the 64-bit register rax, then the following x86 code will set register xmm0 to the corresponding f64:
bt bx, ax # assumes bx initialized to 0x1919 elsewhere
jc self_tagged
mov xmm0,[rax+offset] # xmm0 $\leftarrow$ value field of float
@@
self_tagged:
ror rax, 4 # rotate right 4 places
@@
The number and type of instructions on the hot path is the same as for the boxing operation.
@@ 5.5. C Implementation
This section explains how the above operations can be implemented in C, as this is a common implementation language, in particular it is used by the Bigloo and Gambit Scheme to C compiler used in the experiments of Sections 3 and 6.
Testing for a specific tag or a pair of tags is fairly straightforward to express in C and follows the same principles as the assembly code, so details are skipped. Testing for a set of tags is more challenging in portable C code. It can be implemented with the following pure C function, which is easily inlined by a C compiler (here testing for the tags 000, 011, and 100):
#define TAG_SET ((1<<0)|(1<<3)|(1<<4)) /* 0x19 for tags 000, 011, and 100 */
inline bool has_tag_0_or_3_or_4(int64_t n) {
return (((uint32_t)1 << (n & 31)) & (~(uint32_t)0/0xff * TAG_SET)) != 0;
The expression n & 31 computes a shift count mod 32 and the bitwise-and therefore tests the bit at index n mod 32 of its second operand. The code uses the compile-time constant ~(uint32_t)0/0xff * TAG_SET that repeats the 8 bit mask TAG_SET 4 times to fill a 32 bit word, giving 0x19191919. This achieves the equivalent of testing the bit at index n mod 8 of TAG_SET.
Using this pure C definition with GCC version 13.2.0 and clang version 18.1.3 on x86-64 results in both cases in the non-destructive approach based on the bit test instruction. The expression n & 31 is optimized out by those compilers because the shift instruction on x86 implicitly does this operation. On older versions of those compilers that would implement this with a destructive shift instruction, the bit test approach can be achieved using the following definition that uses an asm statement:
inline bool has_tag_0_or_3_or_4(int64_t n) {
@@
Boxing and unboxing operations need to convert between a float value and its bit representation. This can be achieved portably with a union type whose fields are of type f64 and i64. Moreover, boxing and unboxing need to rotate the bit representation of the float. Although C does not provide an operator for bit rotation, both GCC and clang recognize an equivalent pair of shifts and generate a single machine rotate instruction.
The implementation of boxing in C for the 3-tag variant is:
@@
#define ROTL(n,s) ((int64_t)(((uint64_t)n << s) | ((uint64_t)n >> (64 - s))))
@@
int64_t result = ROTL(((union di)f).i, 4);
if (has_tag_0_or_3_or_4(result)) return result;
return heap_allocate_float(f);
The implementation of unboxing for the 3-tag variant is:
@@
int64_t result = ROTL(o, 60);
if (has_tag_0_or_3_or_4(result)) return ((union di)result).d;
return value_of_heap_allocated_float(o);
6. Experiments of Self-Tagging using the Exponent Bits
The experimental setup from Section 3 is replicated to test float self-tagging using high bits of the exponent. Three variants are compared to the original version of Bigloo that implements standard tagged objects.
  • 3-tag self-tagging uses the tags 000, 011 and 100, thus self-tagging floats in the green, blue, and orange ranges from Figure 6. The remaining floats are represented by generic pointers with a header type, and thus are heap-allocated.
  • 2-tag self-tagging tests the impact of expending the tag 000 for self-tagging. It only uses the tags 011 and 100 for float self-tagging and reclaims the tag 001 to represent all remaining floats as tagged pointers. The floats $\pm$0.0 are preallocated and the float boxing operation contains an explicit zero check to return the preallocated zeros.
  • 1-tag self-tagging tests the 1-tag variant where the tag 001 is used for self-tagged floats.
Figure 8 shows the memory allocations of these variants. All three variants allocate fewer floats on all float-intensive benchmarks. In fact, all variants have been verified to allocate no float at all (see Figure 6) with the only exceptions being the preallocated $\pm$0.0 by the 2-tag variant. The remaining heap usage is for non-float objects. Similarly to the previous experiment (Figure 5(a)), memory usage of other benchmarks remains mostly unaffected.
Figure 8. Heap allocation comparison between the three variants of self-tagging on exponent bits. The 3-tag variant uses the tags 000, 011 and 100. The 2-tag variant uses the tags 011 and 100, but preallocates 0.0 and -0.0. The 1-tag variant uses 3 bits of the biased exponent (see Figure 6). The y-axis shows heap allocations relative to the original Bigloo version, which uses tagged pointers for all floats. 1.0 indicates no change, while 0.0 indicates that all allocations were removed.
Figure 9 shows faster execution times for all three variants of self-tagging. All float-intensive benchmarks execute faster than the original Bigloo implementation, with the exception of fft on the 2-tag variant. In particular, the mbrot and nucleic benchmarks are both faster, confirming that self-tagging on exponent bits eliminates the branch misprediction penalty incurred by self-tagging on mantissa low bits (see Section 3). Despite small individual variations, other benchmarks remain generally unaffected.
On the float-intensive benchmarks, execution time with the 3-tag variant, which makes use of the tag 000, is on average 0.41$\times$ that of the original Bigloo version, slightly more than two times faster. This is significantly lower than the 2-tag variant (0.48$\times$) which avoids the tag 000. In particular, the benchmark fft, which predominantly uses the values $\pm$0.0, is slower in the 2-tag variant despite zeros being preallocated. The 1-tag variant (0.42$\times$) is similar to 3-tag on float benchmarks, and both variants display no performance change on non-float benchmarks (1.01$\times$ for 1-tag, and 1.03$\times$ for 3-tag). Due to the 1-tag variant not using 000, and leaving more tags available, it is a more versatile candidate and is thus used as the default variant for the rest of this paper.
Figure 9. Execution time comparison between the three variants of self-tagging on exponent bits. The 3-tag variant uses the tags 000, 011 and 100. The 2-tag variant uses the tags 011 and 100, but preallocates 0.0 and -0.0. The 1-tag variant uses 3 bits of the biased exponent (see Figure 6). The y-axis shows execution time relative to the original version. 1.0 indicates no change, lower means faster execution than the original Bigloo version, which uses tagged pointers.
6.1. Effect on Garbage Collection
Execution time (s)
fft Refer to caption
fibfp Refer to caption
mbrot Refer to caption
nucleic Refer to caption
pnpoly* Refer to caption
ray Refer to caption
simplex Refer to caption
sum1 Refer to caption
sumfp Refer to caption
Preallocated Heap Data Size
Figure 10. Execution time of float-intensive R7RS benchmarks with extra live data allocated on heap before execution. Benchmarks are executed in unsafe mode (without typechecks) to specifically gauge the impact on garbage collection. The x-axis shows the size of preallocate data in bytes on a logarithmic scale. An extra point is added at $x=0$ for execution time without preallocated data. The y-axis represents execution time in seconds. Dashed lines represent systems that do not implement self-tagging (original Bigloo, original Gambit, and Chez Scheme). Solid lines are Bigloo and Gambit with self-tagging (1-tag).
*pnpoly’s execution time with Bigloo (orig) is about 230 seconds, which is cropped out for readability.
Float self-tagging improves performance by allocating fewer floats on the heap, thus relieving pressure on the garbage collector. Hence, the impact of self-tagging is more pronounced for programs that spend more time doing memory management tasks. This section presents an experiment to measure execution time on float-intensive R7RS benchmarks when the heap contains live data, thus putting more pressure on the garbage collector.
The R7RS benchmarks suite is modified to allocate a vector before the execution of each benchmark. The size of the allocated vector is given as a parameter of the benchmark. Each benchmark is then executed normally with vectors of size ranging from $10^{4}$ to $10^{8}$, which correspond to about 80 kB and 800 MB of memory respectively. An execution is also done with no preallocated vector. To specifically measure the impact of self-tagging on garbage collection, benchmarks are executed in unsafe mode, which allows unsafe optimizations such as the removal of safety typechecks.
For this experiment the float self-tagging approach has been implemented in the Gambit 4.9.6 (Marc Feeley, 2025) optimizing Scheme compiler. The benchmarks are executed with Bigloo and Gambit, with and without self-tagging (1-tag), and also the Chez Scheme 10.1 (R. Kent Dybvig, 2025) optimizing Scheme compiler without self-tagging. These compilers are chosen due to their status as state-of-the-art Scheme implementations and the fact they employ distinct garbage collection techniques. Chez Scheme uses a generational garbage collector. Gambit uses a stop-and-copy garbage collector with bump allocation. Bigloo uses the Boehm-Demers-Weiser Garbage Collector (Boehm and Weiser, 1988).
Figure 10 compares execution times of these Scheme implementations. With a nearly empty heap, and thus little strain on the garbage collector, the bump allocator of Gambit without self-tagging generally outperforms all other implementations since it does not incur the cost, albeit small, of encoding and decoding self-tagged floats. However, once the heap is preloaded with 1 to 10 MB (which is roughly the size of the data cache) the strain on all garbage collectors starts to outweigh the cost of encoding and decoding self-tagged floats, thus Bigloo and Gambit with self-tagging hereafter outperform all implementations without self-tagging.
The threshold of 1 to 10 MB of preallocated data beyond which self-tagging increasingly outperforms other implementations suggests that experiments in other sections are in fact worst-case scenarios. On real life programs, where it is reasonable to expect some persistent data in the heap, self-tagging can be expected to offer a more substantial improvement.
6.2. Comparison to NaN-Boxing and NuN-Boxing
Bigloo also contains implementations of NaN-boxing and NuN-boxing which can be enabled through a configuration option.
Figure 11 compares Bigloo’s original representation (tagged pointers) to NaN-boxing, NuN-boxing, and 1-tag float self-tagging. On float-intensive benchmarks, the average execution time with self-tagging (0.42$\times$ of original version) is higher than with NaN-boxing (0.30$\times$) and NuN-boxing (0.34$\times$). On non-float benchmarks the average relative execution time of self-tagging (1.01$\times$), NaN-boxing (1.03$\times$), NuN-boxing (1.03$\times$) all remain similar to the original version. However, on individual benchmarks the performance of self-tagging remains generally unchanged while significant variations are observed with NaN/NuN-boxing.
Figure 11. Execution time comparison between Bigloo with 1-tag self-tagging, Bigloo with NuN-boxing, and Bigloo with NaN-boxing. The y-axis shows execution time relative to the original Bigloo version, which uses tagged pointers. 1.0 indicates no change, lower means faster execution than the original version.
6.3. Application to JavaScript
This section experiments with the application of float self-tagging in JavaScript, which is a prime candidate for self-tagging because it has no integer type. Its specification defines numbers as double-precision floats (International, 2015). To isolate the impact of the representation of floats, four versions of the Hopc JavaScript compiler are tested, where each version differs only by its representation of floats, which is either tagged pointers, NaN-boxing, NuN-boxing, or self-tagging.
Hopc is an ahead-of-time JavaScript compiler that uses Scheme as its compilation backend and relies on the underlying Scheme compiler for optimizations (Serrano, 2021). This permits reusing Bigloo’s runtime to test float self-tagging in the context of JavaScript.
Figure 12 presents the results of porting Bigloo’s NaN-boxing, NuN-boxing, and self-tagging implementations to JavaScript. Performance is measured using the Hop benchmark suite, which gathers benchmarks from the popular JetStream, Octane and Sunspider JavaScript benchmark suites (Serrano, 2024). Once again, it is useful to categorize benchmarks between float-intensive benchmarks (almabench, n-body, navier-stokes, and raytrace), and benchmarks that use few or no floats.
On the float-intensive benchmarks, the relative execution times of Hopc NaN-boxing (0.26$\times$), NuN-boxing (0.26$\times$), and self-tagging (0.28$\times$) implementations are comparable, roughly 4 times faster than tagged pointers. Execution time with NaN-boxing and NuN-boxing is faster on almabench and n-body, but slower on raytrace and navier-stokes. However, on non-float benchmarks the self-tagging implementation (1.03$\times$) generally fares better due to its lower overhead than NaN-boxing (1.12$\times$) and NuN-boxing (1.16$\times$) when handling non-float objects, by up to a factor of 2.07$\times$ for crypto-sha1.
Figure 12. Execution time comparison between Hopc with 1-tag self-tagging, Hopc with NuN-boxing, and Hopc with NaN-boxing. The y-axis shows execution time relative to the original Hopc version, which uses tagged pointers. 1.0 indicates no change, lower means faster execution than the original version.
4 most 4 most
signif. signif.
expo. expo. Value
bits bits + 1 range
0.0
$0$ 0001 1.4e-45 .. 3.9e-34
0001 0010 3.9e-34 .. 2.5e-29
0010 0011 2.5e-29 .. 1.7e-24
0011 0100 1.7e-24 .. 1.1e-19
0100 0101 1.1e-19 .. 7.1e-15
0101 0110 7.1e-15 .. 4.7e-10
0110 0111 4.7e-10 .. 3.1e-5
0111 1000 3.1e-5 .. 2
1000 1001 2 .. 1.3e5
1001 1010 1.3e5 .. 8.6e9
1010 1011 8.6e9 .. 5.6e14
1011 1100 5.6e14 .. 3.7e19
1100 1101 3.7e19 .. 2.4e24
1101 1110 2.4e24 .. 1.6e29
1110 $1$ 1.6e29 .. 1.0e34
1.0e34 .. 3.4e38
$1$ $0$ Infinity/NaN
Figure 13. The positive value ranges captured by each combination of the 4 most significant exponent bits of the IEEE754 32-bit representation. The rows coloured in bold pink are covered by the 1-tag variant (tag 00) and the rows coloured in light pink are covered by the 2-tag variant (tags 00 and 11).
@@ 7. Self-Tagging on 32-bit Architectures
Although previous sections only considered double-precision floats on 64-bit architectures, self-tagging of single-precision floats is a natural adaptation for 32-bit architectures. Indeed, the core idea, which is to superimpose a tag to a sequence of bits likely to appear in practice, can be used for any machine word size. A popular object representation on 32-bit architectures is to use 2-bit tags (4 possible tags) and heap-allocated objects aligned to 32-bit words.
Figure 13 shows the ranges covered by each combination of the top 4 bits of the exponent, which is 8 bits wide in the IEEE754 32-bit representation.
The 1-tag approach described previously can be adapted to the narrower exponent of the IEEE754 32-bit representation field: the self-tagged encoding of a float whose IEEE754 32-bit representation is the integer $n$ can be computed as $(n\oplus((1+2\times tag)<<27))<<_{rot}4$, where $\oplus$ is addition modulo $2^{32}$ and $<<_{rot}$ is the left rotation operator. If the two lowest bits are equal to $tag$ then the float is self-tagged. This will cover all float values in the ranges: $0.0..3.9\times 10^{-34}\cup 3.1\times 10^{-5}..1.3\times 10^{5}\cup 1.0\times 10^{34}..Infinity/Nan$ (bold pink rows of Figure 13 using $tag$=00).
The coverage can be doubled by using an adaptation of the 2-tag variant described in Section 4.4: $0.0..2.5\times 10^{-29}\cup 4.7\times 10^{-10}..8.6\times 10^{9}\cup 1.6\times 10^{29}..Infinity/Nan$ (light pink rows of Figure 13 using $tag_{1}$=00 and $tag_{2}$=11). With this variant, it would be reasonable to reserve one tag for small signed integers in the range $-2^{29}..2^{29}-1$ (-536870912 .. 536870911), and the remaining tag for heap-allocated objects, including the floats that can’t be self-tagged.
@@ 8. Related Work
Handling floats is a long-standing issue in dynamic language implementation. In the last decades, little progress has been made improving the encoding of floats, with most implementations either using heap-allocated floats or suffering from the overhead of NaN-boxing on pointers. Thus, more general strategies that use data flow analysis (Henglein, 1992; Serrano and Feeley, 1996; Petersen and Glew, 2012; Lindahl and Sagonas, 2003) and partial evaluation (Würthinger et al., 2017) have been developed to find locations where the type of a value is known across its lifecycle. This allows the generation of specialized code that safely handles fully unboxed, untagged values. Such strategies are not specific to unboxing floats. Rather, they tackle the more general problem of inferring types in dynamic languages, which allows dropping type information in either tagged pointers or NaN-boxed pointers. While more limited in scope, self-tagging solves the problem of heap-allocated floats with a new encoding that involves no program analysis.
@@
Self-tagging is straightforward to add to compilers that already use object tagging. Therefore, implementations that represent floats as tagged pointers could benefit from it with minimal implementation effort. Such popular implementations include CPython (Gross, 2023) and Google’s V8 (Sheludko and Solanes, 2020). Compilers that use NaN/NuN-boxing, such as Mozilla’s SpiderMonkey (Foundation, 2024) and Apples’s JavaScriptCore (Inc., 2024) can also benefit from the lower impact that self-tagging has on the performance of non-float types.
@@
Due to the shortcomings of tagged pointers and NaN-boxing, additional strategies are used to avoid boxing. A straightforward approach is to provide homogeneous data structures such as TypedArray in JavaScript (International, 2015) and numpy’s float64 arrays in Python (Harris et al., 2020). Type homogeneity then allows unboxing values within the collection. More generally, compilers can detect data homogeneity at run time and use context-dependent storage strategies to store data in a way that prevents pointer chasing (Bolz et al., 2013). Since self-tagging unboxes most floats in practice, it effectively prevents such pointer chasing in the case of floats without static or run time analysis.
Even for general-purpose languages not specifically designed for numerical computation, ensuring reasonably fast floating-point operations is a major concern. All modern production-quality implementations surveyed for this paper devote significant effort to optimizing them. Unfortunately the academic literature is scarce when it comes to methods and techniques used by these efficient implementations.
In many cases, the source code itself has to be studied to understand the implementation and compare it to self-tagging. The NuN-boxing implementation used in experiments in this paper is based on that of JavaScriptCore [^2] https://github.com/WebKit/WebKit/commit/74c9b9bbc2c7527144ac366c3f62f84aaa1a8a4e. It uses the pattern 0x0000 in the highest 16-bits to encode pointers and 0xffff to encode integers. The other patterns are used for floats. The actual types of heap-values is either encoded in the three lowest bits of the pointer itself or in a header word. For instance, as strings are frequent data type, the pattern 110 is used to encode these values and checking that a value is a string requires to check the lowest three bits and the 16 highest bits. NuN-boxing has the benefits over self-tagging that it encodes all floats uniformly and does not reserve tags for them. Its drawbacks are that it negatively impacts the type checking and accesses of all objects, it can represent integers with at most 48 bits (usually just 32 bits for practical reasons), and it can’t be used on 32-bit architectures.
The technique closest to self-tagging discovered among surveyed implementations are those of CRuby [^3] https://github.com/ruby/ruby/commit/b3b5e626ad69bf22be3228f847f94e1b68f40888 and OpenSmallTalk [^4] https://clementbera.wordpress.com/2018/11/09/64-bits-immediate-floats. Both implementations use a form of self-tagging with a rotation of the mantissa close to the encoding of Section 4.3. However, bit patterns captured by both implementations exclude ±$\pm 0.0$, which is treated as a special case. This imposes complex sequences of instructions for encoding and decoding that can hardly be inlined in the generated code and risks increasing the number of branch misses as demonstrated in Section 3.4.
To the best of the authors’ knowledge, this is the first peer-reviewed paper exposing the principles of self-tagging and evaluating its benefits, drawbacks, and possible variants.